# Supernova-induced binary-interaction-powered supernovae: a model for SN2022jli

### v1

@@ 3.1 Accretion dynamics
In the thermal feedback model (Figure 2), a low-density cavity is formed behind the NS. It is roughly spherical around periastron (panel (1)), but elongates into a large cone at and after apastron (panels (3)–(4)). The cavity severely quenches the accretion rate, by 4similar-toabsent4$\sim$ 4∼ 45555 orders of magnitude, down to around the Eddington limit (Figure 4, light green curve). The effect of such “negative feedback” has been discussed in depth in the context of common-envelope evolution (e.g. Soker, 2016; Grichener et al., 2021; Hillel et al., 2022).
In the bipolar feedback model (Figure 3), a low-density cavity is created in the vertical direction, while the density on the orbital plane is sustained high. The accretion rate is similarly quenched compared to the no feedback model, but only by an order of magnitude around periastron. Around apastron, the accretion is almost completely shut off similar to the thermal feedback model. As a result, the fluctuation amplitude of the mass accretion rate reaches 3similar-toabsent3$\sim$ 3∼ 36666 orders of magnitude. The peak accretion rates steadily decline for each periastron passage, at a rate similar to the no feedback model.
@@
Orbital eccentricity has a stronger influence on the accretion rate. As we can see in Figure 6, the peak accretion rates at periastron clearly correlate with eccentricity, where the $e=0.7$ model has almost an order of magnitude higher accretion rate than the $e=0.5$ model. Interestingly, the $e=0.3$ model seems to lose the clear periodicity and instead shows large fluctuations that are less obviously related to the orbital phase. Note that there is a two-fold effect of varying eccentricity in our setup. Since we fix the orbital period to $P_{\mathrm{orb}}=12.5~{}\mathrm{d}$, the eccentricity determines the periastron distance and hence the pre-SN orbital separation. Higher eccentricity leads to more energy injection through ejecta-companion interaction (for a given explosion energy), as the solid angle subtended by the companion is larger. Therefore, the higher eccentricity models lead to larger inflated radii with more mass in the inflated envelope. The the other effect of higher eccentricity is that with shorter periastron distances, NS will orbit in deeper parts of the envelope. Even if the envelope profiles were the same, the higher eccentricity models would have higher mass accretion rates because the NS reaches down to the deeper, higher density layers.
@@
We also compare the mass ejected from the companion star in Figure 7. To compute the companion mass, we integrate the mass in all bound cells, where a cell is marked as bound when $\phi_{\mathrm{gas}}+v^{2}/2<0$. Here, $\phi_{\mathrm{gas}}$ is the gravitational potential of the gas that does not include the contribution from the NS particle, and $v\equiv|\bm{v}|$ is the cell velocity in the frame of the companion. We see that straight after the SN, $10^{-3}$$10^{-2}~{}M_{\odot}$ of mass is ejected from the companion due to the SN-heating. While only $2\times 10^{-3}~{}M_{\odot}$ is ejected in the $e=0.3$ model (pink curve), $1.2\times 10^{-2}~{}M_{\odot}$ is ejected in the $e=0.7$ model (red curve) due to the larger energy injection via ejecta-companion interaction. The curves with the same orbital parameters but different feedback methods (grey, dark green and light green curves) show almost identical initial mass loss (Figure 4). This confirms that the initial jump is purely due to SN-heating and not because of subsequent energy injection from the feedback. After the second periastron passage, the ejected mass show small deviations between the feedback and no-feedback model, indicating that accretion feedback is causing more mass ejection from the system. Given that the total mass accretion over the duration of our simulations is $\Delta M_{\mathrm{NS}}\lesssim 10^{-4}~{}M_{\odot}$, we can safely say that this mass is ejected from the system and not accreted onto the NS. In the $e=0.7$ model, we can see more significant mass ejection episodes around each periastron passage. The mass ejected per periastron decreases each orbit, with $\sim 10^{-3}~{}M_{\odot}$ ejected in the first passage while only a few $\sim 10^{-4}~{}M_{\odot}$ is ejected in the later orbits.
Figure 7: Time evolution of the mass ejected from the donor star. We compare all models with $M_{2}=5~{}M_{\odot}$. We treat model M05e05b as the “Baseline” model and colour variations are the same as in Figures 4 and 6.
@@ 3.2 Light curve
Figure 8 displays our synthetic light curves based on the hydrodynamic simulations. We show the M05e07b model as it has the highest accretion rate among our models with feedback, but all other models show qualitatively similar behaviour. The grey curve shows the accretion luminosity from the NS particle, which is directly proportional to the mass accretion rate (Eq. (8)). Due to the strong phase-dependence of the accretion, the energy injection into the system can be considered as an almost impulsive explosion at each periastron passage. The resulting light curves have large variations depending on the viewing direction, reflecting the complex geometry of the source binary.
Right after the SN explosion, there is almost no viewing direction dependence. This is because initially, the companion envelope that reprocesses the injected energy expands in an almost spherical manner due to our SN-heating method. After the second periastron passage, the light curves start to significantly deviate from each other due to the envelope being distorted by the orbital interaction. During the first few orbits, the peak luminosities are highest in the $-x$ (pink curve) and $+y$ (blue curve) directions. These angles respectively view the binary from the equatorial plane in the direction of periastron ($-x$) and the direction perpendicular to that on the post-periastron side ($+y$). From this side of the orbit, the NS is in front of the companion star when the accretion rate (and therefore the feedback) is high. On the other side of the orbit ($+x$ and $-y$ directions), the NS is behind the companion when the accretion rate is high, so the radiation is damped by the optically thick inflated envelope before it reaches the observer.
Each successive peak generally gets dimmer and dimmer for two reasons. One is that the accretion rate declines every periastron due to the decline in envelope density. Another effect is the build-up of equatorial obscuring material due to multiple mass ejections at periastron passage. Such mass ejections only occur along the orbital plane, and conversely in the poleward direction (±$\pm z$), the material gets more and more evacuated by the bipolar outflows. This leads to a rise in the peak luminosities as we can see in the purple curve.
@@
All light curves show some level of undulation, where the amplitude strongly depends on the viewing direction. Generally speaking, both the peak luminosity and undulation amplitude is smaller when viewed from the orbital plane compared to the poles. When viewed from intermediate viewing angles (dashed curves), the light curve is generally in between the pole-on light curve (±$\pm z$) and the equatorial light curve of the given azimuthal angle $\varphi$.
Overall, our light curves resemble the key features of the SN2022jli light curve. After $\gtrsim 50~{}\mathrm{d}$, the light curve reaches a state where it steadily declines with some periodic undulations. This is likely the most important characteristic of SN2022jli that sets it apart from other known transients. Typical SNe show declines but without periodic undulations. Some mass-transferring binaries, such as X-ray binaries, show periodic modulations without a long-term decline. The combination of both the decline and undulation is naturally achieved in our model. Quantitatively, the bolometric luminosity is in the same ballpark as that of SN2022jli and so is the decline rate. The undulation amplitude is also consistent if viewed from close to the orbital plane. On the other hand, there are some key features that we do not reproduce, including the first $\lesssim 50~{}\mathrm{d}$ and the drop at $\sim 250~{}\mathrm{d}$. We discuss these points in Section 4.
3.3 Photospheric evolution
As we show in Figure 9, the effective temperature steadily decreases while the photospheric radius gradually increases over time in our models. There are periodic signatures both in the temperature and radius, which become weaker for the edge-on viewing directions. The rise in photospheric radius seems to stall around the fifth periastron, and may be starting to decline towards the end of our simulation. This plateau-like behaviour is due to the opacity having a steep drop around the recombination temperature $\sim 10^{4}~{}\mathrm{K}$. Both temperature and radius roughly have the same order of magnitude as observed in SN2022jli ($T_{\mathrm{eff}}\sim 10^{4}~{}\mathrm{K},R_{\mathrm{eff}}\sim 10^{15}~{}\mathrm{cm}$). The photospheric radius is still well contained within our computational domain ($8\times 10^{15}~{}\mathrm{cm}$), validating our choice of the outer boundary radius. While we do not see the overall blue-ward evolution seen in SN2022jli, the model reproduces the correct phase evolution where the colour is bluer at undulation peaks (Cartier et al., 2024).
The complex structure of the photosphere can be viewed more clearly in Figures 10 & 11. Here, we plot 1D profiles along various radial rays at given snapshots. The curves are colour-coded based on the latitudinal angle, with darker colours towards the pole and lighter colours towards the equator. We can clearly see the surface layers of the star have been significantly inflated from the initial profile (grey dashed curve), extending out to the location of the NS and beyond.
At periastron (Figure 10), we can see the low-$\theta$ curves have much lower density compared to the high-$\theta$ curves around the NS (vertical line), which is because of the bipolar cavity blown out by the feedback. The profile interior to the location of the NS is roughly spherical with low velocities, indicating it is roughly in hydrostatic equilibrium. The exterior is extremely asymmetrical, showing $\lesssim 5$ orders of magnitude variations in density depending on the direction. A similar trend can be seen at apastron (Figure 11), with now a larger hydrostatic interior region and strongly perturbed exterior. Generally, the equatorial profiles seem to have higher densities compared to the polar profiles, showing that the outflow is equatorially concentrated.
@@ 3.3 Photospheric evolution
Figure 10: Profiles of density, pressure, temperature, opacity and radial velocity as a function of distance from the companion core. Each curve corresponds to the density profile along a specific radial ray and are coloured according to the latitudinal angle. The brown vertical lines mark the location of the NS. We choose a snapshot from model M05e07b at $t=37.5~{}\mathrm{d}$, which is the third periastron. The grey dashed curves show the initial profiles.
Figure 11: Same as Figure 10 but for $t=43.75~{}\mathrm{d}$; the fourth apastron.
@@ 3.4 Undulation features
We now compare the shape of the undulations between our model and observations. We focus on the light curve for the $+x$ viewing direction in Figure 8 (red curve) as it has the smallest undulation amplitude. To extract the undulations, we first fit a sixth-order polynomial curve to the light curve between 50 and 100 d to obtain a baseline. We plot the difference between the light curve and the baseline against the undulation phase in Figure 12. The observed undulation light curve is computed in the same way and overplotted. Overall, the model is in good agreement with the observed shape. There is a sharp rise followed by a longer decline. Each cycle has slightly different shapes both in the model and observations, partly due to the ambiguity in defining the baseline, but also due to the stochastic nature of the undulations.
Figure 12: Undulation light curve folded over undulation phase. Grey points are the observed data of SN2022jli (Chen et al., 2024). Red solid curves are the folded light curves of the $+x$ model in Figure 8. Both are shifted so that the minima are located at the phase origin.
While our edge-on light curve models closely agree with the small amplitude undulations in SN2022jli, we also predict large amplitude oscillations if a similar event was viewed from higher-latitude viewing directions. Recently, another rare kind of stripped-envelope SN was discovered which re-emerged as a type IIn SN, $\sim 3$ yr after the first event (SN2022mop; Brennan et al., 2025). In both the tail of the first event and the rise towards the second event, several periodic outbursts ($\sim 1$ mag) were detected with $\sim 27~{}\mathrm{d}$ intervals. The shape and amplitude of the outbursts loosely resemble the saw-tooth like morphology that we find in our hydrodynamic models observed from the pole (purple curves in Fig 8). It may be that the first event of SN2022mop was an event similar to SN2022jli, but with different orbital periods and viewed from different angles. Brennan et al. (2025) report that SN2009ip also has periodic outbursts with amplitudes and intervals similar to SN2022mop.

### v2

@@ 3.1 Accretion dynamics
In the thermal feedback model (Figure 2), a low-density cavity is formed behind the NS. It is roughly spherical around periastron (panel (1)), but elongates into a large cone at and after apastron (panels (3)–(4)). The cavity severely quenches the accretion rate, by $\sim 4$55 orders of magnitude, down to around the Eddington limit (Figure 4, light green curve). The effect of such “negative feedback” has been discussed in depth in the context of common-envelope evolution (e.g. Soker, 2016; Grichener et al., 2021; Hillel et al., 2022).
In the bipolar feedback model (Figure 3), a low-density cavity is created in the vertical direction, while the density on the orbital plane is sustained high. The accretion rate is similarly quenched compared to the no feedback model, but only by an order of magnitude around periastron. Around apastron, the accretion is almost completely shut off similar to the thermal feedback model. As a result, the fluctuation amplitude of the mass accretion rate reaches $\sim 3$66 orders of magnitude. The peak accretion rates steadily decline for each periastron passage, at a rate similar to the no feedback model.
@@
Orbital eccentricity has a stronger influence on the accretion rate. As we can see in Figure 6, the peak accretion rates at periastron clearly correlate with eccentricity, where the $e=0.7$ model has almost an order of magnitude higher accretion rate than the $e=0.5$ model. Interestingly, the $e=0.3$ model seems to lose the clear periodicity and instead shows large fluctuations that are less obviously related to the orbital phase. Note that there is a two-fold effect of varying eccentricity in our setup. Since we fix the orbital period to $P_{\mathrm{orb}}=12.5~\mathrm{d}$, the eccentricity determines the periastron distance and hence the pre-SN orbital separation. Higher eccentricity leads to more energy injection through ejecta-companion interaction (for a given explosion energy), as the solid angle subtended by the companion is larger. Therefore, the higher eccentricity models lead to larger inflated radii with more mass in the inflated envelope. The other effect of higher eccentricity is that with shorter periastron distances, the NS will orbit in deeper parts of the envelope. Even if the envelope profiles were the same, the higher eccentricity models would have higher mass accretion rates because the NS reaches down to the deeper, higher density layers.
@@
Figure 7 compares the optimistic model and the realistic model for the mass accretion rate. In a short test simulation (purple curve), we found that the realistic model shows $\gtrsim 15$ times lower accretion rates at periastron compared to the optimistic model with the same orbital parameters (red curve). The feedback luminosity is therefore far too weak to match that of SN2022jli. A model with higher eccentricity $e=0.8$ (black curve) roughly compensates for this loss, showing very similar mass accretion rate evolution as the $e=0.7$ model with optimistic mass accretion. Therefore, in terms of mass accretion rate evolution, there exists a degeneracy between the mass accretion efficiency ($p$) and the orbital eccentricity.
Figure 7: Same as Figure 4 but comparing the optimistic (model M05e07b; red) vs realistic (model M05e07b_r; purple, M05e08b_r; dark brown) models for the mass accretion rate.
We also compare the mass ejected from the companion star in Figure 8. To compute the companion mass, we integrate the mass in all bound cells, where a cell is marked as bound when $\phi_{\mathrm{gas}}+v^{2}/2<0$. Here, $\phi_{\mathrm{gas}}$ is the gravitational potential of the gas that does not include the contribution from the NS particle, and $v\equiv|\bm{v}|$ is the cell velocity in the frame of the companion. We see that straight after the SN, $10^{-3}$$10^{-2}~M_{\odot}$ of mass is ejected from the companion due to the SN-heating. While only $2\times 10^{-3}~M_{\odot}$ is ejected in the $e=0.3$ model (pink curve), $1.2\times 10^{-2}~M_{\odot}$ is ejected in the $e=0.7$ model (red curve) and $3.5\times 10^{-2}~M_{\odot}$ in the $e=0.8$ model (dark brown curve) due to the larger energy injection via ejecta-companion interaction. The curves with the same orbital parameters but different feedback methods (grey, dark green and light green curves) show almost identical initial mass loss (Figure 4). This confirms that the initial jump is purely due to SN-heating and not because of subsequent energy injection from the feedback. After the second periastron passage, the ejected mass show small deviations between the feedback and no-feedback model, indicating that accretion feedback is causing more mass ejection from the system. Given that the total mass accretion over the duration of our simulations is $\Delta M_{\mathrm{NS}}\lesssim 10^{-4}~M_{\odot}$, we can safely say that this mass is ejected from the system and not accreted onto the NS. There is hardly any mass lost in the no-feedback model, indicating that tidal stripping is negligible on these timescales. In the high eccentricity models ($e=0.7,0.8$), we can see more significant mass ejection episodes around each periastron passage. These results indicate that the depth of energy injection is important in determining the ejected mass as well as the amount.
Figure 8: Time evolution of the mass ejected from the donor star. We compare all models with $M_{2}=5~M_{\odot}$. We treat model M05e05b as the “Baseline” model and colour variations are the same as in Figures 4, 6 and 7.
@@ 3.2 Light curve
Figure 9 displays our synthetic light curves based on the M05e07b model, which uses the optimistic mass accretion rate and an orbital eccentricity $e=0.7$. The grey curve shows the accretion luminosity from the NS particle, which is directly proportional to the mass accretion rate (Eq. (11)). Due to the strong phase-dependence of the accretion, the energy injection into the system can be considered as an almost impulsive explosion at each periastron passage. The resulting light curves have large variations depending on the viewing direction, reflecting the complex geometry of the source binary.
Very roughly, the overall light curve shows a steady decline after the first 1122 orbits. On top of that, there are strong periodic undulations peaking at periastron. Each successive peak generally gets dimmer and dimmer for two reasons. One is that the accretion rate declines every periastron due to the decline in envelope density. Another effect is the build-up of equatorial obscuring material due to multiple mass ejections at periastron passage. Such mass ejections only occur along the orbital plane, and conversely in the poleward direction ($\pm z$), the material gets more and more evacuated by the bipolar outflows. This leads to a rise in the peak luminosities as we can see in the purple curve.
Right after the SN explosion, there is almost no viewing direction dependence. This is because initially, the companion envelope that reprocesses the injected energy expands in an almost spherical manner due to our SN-heating method. After the second periastron passage, the light curves start to significantly deviate from each other due to the envelope being distorted by the orbital interaction. During the first few orbits, the peak luminosities are highest in the x-x (pink curve) and +y+y (blue curve) directions. These angles respectively view the binary from the equatorial plane in the direction of periastron (x-x) and the direction perpendicular to that on the post-periastron side (+y+y). From this side of the orbit, the NS is in front of the companion star when the accretion rate (and therefore the feedback) is high. On the other side of the orbit (+x+x and y-y directions), the NS is behind the companion when the accretion rate is high, so the radiation is damped by the optically thick inflated envelope before it reaches the observer.
@@
All light curves show some level of undulation, where the amplitude strongly depends on the viewing direction. Generally speaking, both the peak luminosity and undulation amplitude is smaller when viewed from the orbital plane compared to the poles. When viewed from intermediate viewing angles (dashed curves), the light curve is generally in between the pole-on light curve ($\pm z$) and the equatorial light curve of the given azimuthal angle $\varphi$.
Figure 10 displays the light curves computed for Model M05e08b_r, which uses the realistic mass accretion model and $e=0.8$. There is a striking difference to that of the M05e07b model light curve. First, the overall light curve slowly rises over a timescale of $\sim 60~\mathrm{d}$ before showing a steady decline. Second, the undulation amplitudes are significantly smaller over all viewing directions. As seen in Figure 7, the mass accretion rate histories are very similar between these models, so the main difference is in the orbital configuration. In the higher eccentricity model, the periastron passage occurs at deeper layers inside the inflated envelope. Therefore, the feedback outflow interacts with more material as it expands, diminishing its energy before being released at the photosphere. The bipolar outflow is also smeared out over a larger solid angle, reducing the viewing direction dependency of the light curve.
Figure 10: Same as Figure 10 but for the realistic mass accretion model.
Overall, our light curves computed with the realistic mass accretion rates resemble the key features of the SN2022jli light curve. After $\gtrsim 60~\mathrm{d}$, the light curve reaches a state where it steadily declines with some periodic undulations. This is likely the most important characteristic of SN2022jli that sets it apart from other known transients. Typical SNe show declines but without periodic undulations. Some mass-transferring binaries, such as X-ray binaries, show periodic modulations without a long-term decline. The combination of both the decline and undulation is naturally achieved in our model. Quantitatively, the bolometric luminosity is in the same ballpark as that of SN2022jli and so is the decline rate. The undulation amplitude depends strongly on the eccentricity and to some degree on the viewing direction, but our $e=0.8$ light curves are in good agreement with SN2022jli. On the other hand, there are some key features that we do not reproduce, including the first $\lesssim 30~\mathrm{d}$ and the drop at $\sim 250~\mathrm{d}$. We discuss these points in Section 4.
3.3 Photospheric evolution
Figures 11 & 12 display the evolution of effective temperature and photospheric radius for models M05e07b and M05e08b_r, respectively. In the lower eccentricity model (M05e07b), the effective temperature steadily decreases while the photospheric radius gradually increases over time in our models. There are periodic signatures both in the temperature and radius, which become weaker for the edge-on viewing directions and/or higher eccentricity. The rise in photospheric radius stalls later on, and may be starting to decline towards the end of our simulation. This plateau-like behaviour is due to the opacity having a steep drop around the recombination temperature $\sim 10^{4}~\mathrm{K}$. Both temperature and radius roughly have the same order of magnitude as observed in SN2022jli ($T_{\mathrm{eff}}\sim 10^{4}~\mathrm{K},R_{\mathrm{eff}}\sim 10^{15}~\mathrm{cm}$). The photospheric radius is still well contained within our computational domain ($8\times 10^{15}~\mathrm{cm}$), validating our choice of the outer boundary radius. While we do not see the overall blue-ward evolution seen in SN2022jli, the model reproduces the correct phase evolution where the colour is bluer at undulation peaks (Cartier et al., 2024).
@@ 3.3 Photospheric evolution
Figure 12: Same as Figure 11 but for the realistic mass accretion model.
The complex structure of the photosphere can be viewed more clearly in Figures 13 & 14. Here, we plot 1D profiles along various radial rays at given snapshots. The curves are colour-coded based on the latitudinal angle, with darker colours towards the pole and lighter colours towards the equator. We can clearly see the surface layers of the star have been significantly inflated from the initial profile (grey dashed curve), extending out to the location of the NS and beyond.
At periastron (Figure 13), we can see the low-$\theta$ (polar) curves have much lower density compared to the high-$\theta$ (equatorial) curves around the NS (vertical line), which is because of the bipolar cavity blown out by the feedback. The profile interior to the location of the NS is roughly spherical with low velocities, indicating it is roughly in hydrostatic equilibrium. The exterior is extremely asymmetrical, showing $\lesssim 5$ orders of magnitude variations in density depending on the direction. A similar trend can be seen at apastron (Figure 14), with now a larger hydrostatic interior region and strongly perturbed exterior. Generally, the equatorial profiles seem to have higher densities compared to the polar profiles, showing that the outflow is equatorially concentrated.
@@
As a visual guide, we overplot the density profile of an example type Ic SN ejecta (red curve). See Appendix B for details of how this was computed. We can see that the bulk of the SN ejecta lie at $r\gtrsim 2\times 10^{4}~R_{\odot}$ and the density of the inner parts of the ejecta are negligible compared to that of the inflated envelope at radii below that. This indicates that the interaction between the inflated envelope and the inner parts of the SN ejecta are weak and will not influence the dynamics in our simulations at least up to the photosphere. Both the inner parts of the SN ejecta and the outer parts of the inflated envelope could be compressed into a thin shell at around $\sim 10^{4}~R_{\odot}$, as they interact. However, the mass in this thin shell is expected to be negligible compared to the rest of the SN ejecta. As we discuss in Section 4.2, the SN ejecta may still be optically thick at these times but could become optically thin later on ($\gtrsim 50~\mathrm{d}$).
Figure 13: Profiles of density, pressure, temperature, opacity and radial velocity as a function of distance from the companion core. Each curve corresponds to the density profile along a specific radial ray and are coloured according to the latitudinal angle. The brown vertical lines mark the location of the NS. We choose a snapshot from model M05e08b_r at $t=50~\mathrm{d}$, which is the fourth periastron. The grey dashed curves show the initial profiles. The overlaid red solid curve is an example ejecta density profile of a type Ic SN.
Figure 14: Same as Figure 13 but for $t=56.25~\mathrm{d}$; the fifth apastron.
@@ 3.4 Undulation features
We now compare the shape of the undulations between our model and observations. We focus on the light curve for the +x+x viewing direction in Figure 9 (red curve). To extract the undulations, we first fit a sixth-order polynomial curve to the light curve between 50 and 100 d to obtain a baseline. We plot the difference between the light curve and the baseline against the undulation phase in Figure 15. The observed undulation light curve is computed in the same way and overplotted. Overall, the model is in good agreement with the observed shape. There is a sharp rise followed by a longer decline. Each cycle has slightly different shapes both in the model and observations, partly due to the ambiguity in defining the baseline, but also due to the stochastic nature of the undulations. Our M05e08b_r model has similar undulation shapes, although the rise is not as sharp as the lower eccentricity model.
Figure 15: Undulation light curve folded over undulation phase for the M05e07b model. Grey points are the observed data of SN2022jli (Chen et al., 2024). Red solid curves are the folded light curves of the +x+x model in Figure 9. Both are shifted so that the minima are located at the phase origin.
While our edge-on light curve models closely agree with the small amplitude undulations in SN2022jli, we also predict large amplitude oscillations if a similar event was viewed from higher-latitude viewing directions and slightly lower eccentricity. Recently, another rare kind of stripped-envelope SN was discovered which re-emerged as a type IIn SN, $\sim 3$ yr after the first event (SN2022mop; Brennan et al., 2025). In both the tail of the first event and the rise towards the second event, several periodic outbursts ($\sim 1$ mag) were detected with $\sim 27~\mathrm{d}$ intervals. The shape and amplitude of the outbursts loosely resemble the saw-tooth like morphology that we find in our hydrodynamic models observed from the pole (purple curves in Fig 9). It may be that the first event of SN2022mop was an event similar to SN2022jli, but with different orbital periods and viewed from different angles or had a lower eccentricity. Brennan et al. (2025) report that SN2009ip also has periodic outbursts with amplitudes and intervals similar to SN2022mop.